recall how to determine the values of the random variable by answering the given problem. Find the values of the ramdon variable Y representing the number of green balls when 2 balls are drawn in succession without replacement from a jar 4 red balls and 5 green balls


1
Expert's answer
2021-03-01T15:38:28-0500

Total balls in the jar: 4+5=9


find the probability that the green balls will be 0, 1 and 2 respectively:


"P(Y = 0) = \frac{4}{9} \cdot \frac{3}{8} = \frac{1}{6}" (both red balls)


"P(Y = 1) = \frac{4}{9} \cdot \frac{5}{8} + \frac{5}{9} \cdot \frac{4}{8} = \frac{5}{9}" (red ball and green ball or green ball and red ball)


"P(Y = 2) = \frac{5}{9} \cdot \frac{4}{8} = \frac{5}{{18}}" (both green balls)


we have the distribution series of the random variable Y:

"\begin{matrix}
Y&0&1&2\\
p&{\frac{1}{6}}&{\frac{5}{9}}&{\frac{5}{{18}}}
\end{matrix}"